The circular plates of a parallel-plate capacitor have a radius of 30 mm. A steady 2.0-A current is charging the initially uncharged capacitor, and the surface charge on the plates is distributed uniformly. Derive an expression for the magnitude of the electric field between the plates as a function of time t where t is in seconds. Express your answer in terms of t
Answer #1

Here,

radius , r = 30 mm = 0.030 m

current , I = 2 A

now , the charge on the capacitor , Q = curent * t

Q = 2 * t

Now , electric field between the plates of capacitor

electric field between the plates of capacitor = Q/(epsilon * Area)

electric field between the plates of capacitor = 2 * t/(8.854 *10^-12 * pi * 0.030^2)

electric field between the plates of capacitor = 7.99 *10^-13 * t N/C

the electric field between the plates of capacitor is 7.99 *10^-13 * t N/C